The data represents the quiz scores of psychology majors in one of their major subjects.


Mean = 22.5 SD = 5.50 N = 60


  1. Find out the number of students with scores equal to and above 30.
  2. How many students have scores equal to and below 30? 
  3. Find out the range of scores of the middle 10 students.
1
Expert's answer
2022-01-06T10:28:24-0500

mean=22.5

Sd= 5.5

N=60

a). Students that score more than 30

P(x"\ge" 30)=(1-P(x <30)

P(x<30)=NORMDIST(30,22.5,5.5,1) =0.913659

1-0.913659

=0.086341

The number above 30 =60*0.086341

Approximately 5 students


b). students who scored less than 30

=NORMDIST (30,22.5,5.5,1);

0.913659 *60

= Approximately 55 Students scored 30 and below.


c).the range of 10 middle student

10/60 =0.1667

The prob is 0.5±0.1667

The probability below is 0.3333

The probability above is 0.6667

The value is given by

NORMINV (0.3333,22.5,5.5)

=20.1305

NORMINV (0.6667,22.5,5.5)

=24.87404

The range is between 20.1305 and 24.87404